OpenStudy (anonymous):

The sum of two numbers is 16. The sum of their squares is a minimum. Determine the number

OpenStudy (anonymous):

WHAT DOES THIS MEAN D:

OpenStudy (aum):

Let the two numbers be: x and (16-x) You need to minimize x^2 + (16-x)^2

OpenStudy (aum):

x^2 + (16-x)^2 is a vertical parabola that opens upward. Minimum will be at the vertex.

OpenStudy (anonymous):

8,8

OpenStudy (anonymous):

its saying the sum of 2 numbers which is x and y so the equation can be like x + y = 16 or y = -x + 16 the sum of two squares can be written as x^2 + y^2 because of the first equation then x^2 + (-x + 16)^2 = x^2 + (-x +16)(-x + 16) then expand the binomial you will get x^2 + x^2 - 16x - 16x + 256 2x^2 - 32x + 256

OpenStudy (anonymous):

x^2 + (16-x)^2 =2x^2 -32x +196 =2(x-8)^2 +132 so min when x=8

OpenStudy (anonymous):

there are 2 ways to find the minimum for that function -b/2a -32/4 = 8

OpenStudy (anonymous):

x^2 + (16-x)^2 =2x^2 -32x +256 (error in calculation) =2(x-8)^2 +192 so min when x=8

OpenStudy (anonymous):

min is 8

OpenStudy (anonymous):

thanks guys! i appreciate your help :D

OpenStudy (anonymous):

yw